# Mean-Field Sampling for Cooperative Multi-Agent Reinforcement Learning

### v1

@@ Acknowledgements.
We gratefully acknowledge insightful discussions with Siva Theja Maguluri, Sarah Liaw, Yiheng Lin, and Yi Wu.
@@ Appendix A Mathematical Background and Additional Remarks
  • Section 5 presents the proof of the Lipschitz continuity between $\widehat{Q}_{k}$ and $Q^{*}$
  • Section 6 presents the TV distance bound
  • Section 7 proves the bound on the optimality gap between the learned policy $\widetilde{\pi}_{k,m}^{\mathrm{est}}$ and the optimal policy $\pi^{*}$
@@
$a_{1},\dots,a_{n}$ $a_{1},\dots,a_{n}\in\mathcal{A}_{l}^{n}$ are the actions of the local agents $1,\dots,n$
$s_{1},\dots,s_{n}$ $s_{1},\dots,s_{n}\in\mathcal{S}_{l}^{n}$ are the states of the local agents $1,\dots,n$
$a$ $a=(a_{g},a_{1},\dots,a_{n})\in\mathcal{A}_{g}\times\mathcal{A}_{l}^{n}$ is the tuple of actions of all agents
$s$ $s=(s_{g},s_{1},\dots,s_{n})\in\mathcal{S}_{g}\times\mathcal{S}_{l}^{n}$ is the tuple of states of all agents
$z_{i}$ $z_{i}=(s_{i},a_{i})\in\mathcal{Z}_{l}$, for $i\in[n]$
$\mu_{k}(\mathcal{Z}_{l})$ $\mu_{k}(\mathcal{Z}_{l})=\{0,1/k,2/k,\dots,1\}^{|\mathcal{Z}_{l}|}$
$\mu(\mathcal{Z}_{l})$ $\mu(\mathcal{Z}_{l})\coloneq\mu_{n}(\mathcal{Z}_{l})\{0,1/n,2/n,\dots,1\}^{|\mathcal{Z}_{l}|}$
$s_{\Delta}$ For $\Delta\subseteq[n]$, and a collection of variables $\{s_{1},\dots,s_{n}\}$, $s_{\Delta}\coloneq\{s_{i}:i\in\Delta\}$
@@
$\widehat{\pi}_{k}^{*}$ $\widehat{\pi}_{k}^{*}$ is the optimal deterministic policy function on a constrained system of $|\Delta|=k$ local agents
$\widetilde{\pi}^{\mathrm{est}}_{k}$ $\widetilde{\pi}^{\mathrm{est}}_{k}$ is the stochastic policy mapping learned with parameter $k$ such that $a\sim\widetilde{\pi}_{k}^{\mathrm{est}}(s)$
$P_{g}(\cdot|s_{g},a_{g})$ $P_{g}(\cdot|s_{g},a_{g})$ is the stochastic transition kernel for the state of the global agent
$P_{l}(\cdot|a_{i},s_{i},s_{g})$ $P_{l}(\cdot|a_{i},s_{i},s_{g})$ is the stochastic transition kernel for the state of any local agent $i\in[n]$
$r_{g}(s_{g},a_{g})$ $r_{g}$ is the global agent’s component of the reward
$r_{l}(s_{i},s_{g},a_{i})$ $r_{l}$ is the component of the reward for local agent $i\in[n]$
$r(s,a)$ $r(s,a)\coloneq r_{[n]}(s,a)=r_{g}(s_{g},a_{g})+\frac{1}{n}\sum_{i\in[n]}r_{l}(s_{i},s_{g},a_{i})$ is the reward of the system
$r_{\Delta}(s,a)$ $r_{\Delta}(s,a)=r_{g}(s_{g},a_{g})+\frac{1}{|\Delta|}\sum_{i\in\Delta}r_{l}(s_{i},s_{g},a_{i})$ is the constrained system’s reward with $|\Delta|=k$ local agents
$\mathcal{T}$ $\mathcal{T}$ is the centralized Bellman operator
$\widehat{\mathcal{T}}_{k}$ $\widehat{\mathcal{T}}_{k}$ is the Bellman operator on a constrained system of $|\Delta|=k$ local agents
@@
Given two metric spaces $(\mathcal{X},d_{\mathcal{X}})$ and $(\mathcal{Y},d_{\mathcal{Y}})$ and a constant $L\in\mathbb{R}_{+}$, a mapping $f:\mathcal{X}\to\mathcal{Y}$ is $L$-Lipschitz continuous if for all $x,y\in\mathcal{X}$, $d_{\mathcal{Y}}(f(x),f(y))\leq L\cdot d_{\mathcal{X}}(x,y)$.
Theorem A.2 (Banach-Caccioppoli fixed point theorem (Banach, 1922)).
@@
Appendix B Notation and Basic Lemmas
@@
Definition B.1 (Bellman Operator $\mathcal{T}$).
@@
Definition B.2 (Adapted Bellman Operator $\widehat{\mathcal{T}}_{k}$).
The adapted Bellman operator updates a smaller $Q$ function (which we denote by $\widehat{Q}_{k}$), for a surrogate system with the global agent and $k\in[n]$ local agents denoted by $\Delta$, using mean-field value iteration and $j\in\Delta$ such that:
$\widehat{\mathcal{T}}_{k}\widehat{Q}_{k}^{t}(s_{g},s_{j},F_{z_{\Delta\setminus j}},a_{j},a_{g}):=r_{\Delta}(s,a)+\gamma\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim P_{g}(\cdot|s_{g},a_{g}),\\ s_{i}^{\prime}\sim P_{l}(\cdot|s_{i},s_{g},a_{i}),\forall i\in\Delta\end{subarray}}\max_{a^{\prime}\in\mathcal{A}}\widehat{Q}_{k}^{t}(s_{g}^{\prime},s_{j}^{\prime},F_{z^{\prime}_{\Delta\setminus j}},a^{\prime}_{j},a^{\prime}_{g})$ (19)
Definition B.3 (Empirical Adapted Bellman Operator $\widehat{\mathcal{T}}_{k,m}$).
The empirical adapted Bellman operator $\widehat{\mathcal{T}}_{k,m}$ empirically estimates the adapted Bellman operator update using mean-field value iteration by drawing $m$ random samples of $s_{g}\sim P_{g}(\cdot|s_{g},a_{g})$ and $s_{i}\sim P_{l}(\cdot|s_{i},s_{g},a_{i})$ for $i\in\Delta$, where for $\ell\in[m]$, the $\ell$’th random sample is given by $s_{g}^{\ell}$ and $s_{\Delta}^{\ell}$, and $j\in\Delta$:
$\widehat{\mathcal{T}}_{k,m}\widehat{Q}_{k,m}^{t}(s_{g},s_{j},F_{z_{\Delta\setminus j}},a_{j},a_{g}):=r_{\Delta}(s,a)+\frac{\gamma}{m}\sum_{\ell\in[m]}\max_{a^{\prime}\in\mathcal{A}}\widehat{Q}_{k,m}^{t}(s_{g}^{\ell},s_{j}^{\ell},F_{z_{\Delta\setminus j}^{\ell}},a_{j}^{\ell},a_{g}^{\ell})$ (20)
Remark B.4.
Since the local agents $1,\dots,n$ are all homogeneous in their state/action spaces, the $\widehat{Q}_{k}$-function only depends on them through their empirical distribution $F_{z_{\Delta}}$. Therefore, throughout the remainder of the paper, we will use $\widehat{Q}_{k}(s_{g},F_{z_{\Delta}},a_{g})\coloneq\widehat{Q}_{k}(s_{g},s_{i},F_{z_{\Delta\setminus i}},a_{g},a_{i})\coloneq\widehat{Q}_{k}(s_{g},s_{\Delta},a_{g},a_{\Delta})$ interchangeably, unless making a remark about the computational complexity of learning each function.
Lemma B.5.
For any $\Delta\subseteq[n]$ such that $|\Delta|=k$, suppose $0\leq r_{\Delta}(s,a)\leq\widetilde{r}$. Then, for all $t\in\mathbb{N}$, $\widehat{Q}_{k}^{t}\leq\frac{\widetilde{r}}{1-\gamma}$.
@@
The proof follows by induction on $t$. The base case follows from $\widehat{Q}_{k}^{0}:=0$. For the induction, note that by the triangle inequality $\|\widehat{Q}_{k}^{t+1}\|_{\infty}\leq\|r_{\Delta}\|_{\infty}+\gamma\|\widehat{Q}_{k}^{t}\|_{\infty}\leq\widetilde{r}+\gamma\frac{\widetilde{r}}{1-\gamma}=\frac{\widetilde{r}}{1-\gamma}$. ∎
Remark B.6.
By the law of large numbers, $\lim_{m\to\infty}\widehat{\mathcal{T}}_{k,m}=\widehat{\mathcal{T}}_{k}$, where the error decays in $O(1/\sqrt{m})$ by the Chernoff bound. Also, $\widehat{\mathcal{T}}_{n}\coloneq\mathcal{T}$. Further, Lemma B.5 is independent of the choice of $k$. Therefore, for $k=n$, this implies an identical bound on $Q^{t}$. An identical argument implies the same bound on $\widehat{Q}_{k,m}^{t}$.
$\mathcal{T}$ satisfies a $\gamma$-contractive property under the infinity norm (Watkins & Dayan, 1992). We similarly show that $\widehat{\mathcal{T}}_{k}$ and $\widehat{\mathcal{T}}_{k,m}$ satisfy a $\gamma$-contractive property under infinity norm in Lemmas B.7 and B.8.
Lemma B.7.
$\widehat{\mathcal{T}}_{k}$ satisfies the $\gamma$-contractive property under infinity norm:
$\|\widehat{\mathcal{T}}_{k}\widehat{Q}_{k}^{\prime}-\widehat{\mathcal{T}}_{k}\widehat{Q}_{k}\|_{\infty}\leq\gamma\|\widehat{Q}_{k}^{\prime}-\widehat{Q}_{k}\|_{\infty}$ (21)
@@
Suppose we apply $\widehat{\mathcal{T}}_{k}$ to $\widehat{Q}_{k}(s_{g},F_{z_{\Delta}},a_{g})$ and $\widehat{Q}^{\prime}_{k}(s_{g},F_{z_{\Delta}},a_{g})$ for $|\Delta|=k$. Then:
$\|\widehat{\mathcal{T}}_{k}\widehat{Q}_{k}^{\prime}-\widehat{\mathcal{T}}_{k}\widehat{Q}_{k}\|_{\infty}$ $=\gamma\max_{\begin{subarray}{c}s_{g}\in\mathcal{S}_{g},\\ a_{g}\in\mathcal{A}_{g},\\ F_{z_{\Delta}}\in\mu_{k}{(\mathcal{Z}_{l})}\end{subarray}}\!\left|\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim P_{g}(\cdot|s_{g},a_{g}),\\ s_{i}^{\prime}\sim P_{l}(\cdot|s_{i},s_{g},a_{i}),\\ \forall i\in\Delta,\\ \end{subarray}}\max_{a^{\prime}\in\mathcal{A}}\widehat{Q}_{k}^{\prime}(s_{g}^{\prime},F_{z_{\Delta}^{\prime}},a_{g}^{\prime})-\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim P_{g}(\cdot|s_{g},a_{g}),\\ s_{i}^{\prime}\sim P_{l}(\cdot|s_{i},s_{g},a_{i}),\\ \forall i\in\Delta^{\prime}\end{subarray}}\max_{a^{\prime}\in\mathcal{A}}\widehat{Q}_{k}(s_{g}^{\prime},F_{z_{\Delta}^{\prime}},a_{g}^{\prime})\right|$
$\leq\gamma\max_{\begin{subarray}{c}s_{g}^{\prime}\in\mathcal{S}_{g},F_{z_{\Delta}}\in\mu_{k}{(\mathcal{Z}_{l})},a^{\prime}\in\mathcal{A}\end{subarray}}\left|\widehat{Q}_{k}^{\prime}(s_{g}^{\prime},F_{z_{\Delta}^{\prime}},a_{g}^{\prime})-\widehat{Q}_{k}(s_{g}^{\prime},F_{z_{\Delta}^{\prime}},a_{g}^{\prime})\right|=\gamma\|\widehat{Q}_{k}^{\prime}-\widehat{Q}_{k}\|_{\infty}$
@@
Lemma B.8.
$\widehat{\mathcal{T}}_{k,m}$ satisfies the $\gamma$-contractive property under infinity norm.
@@
Similarly to Lemma B.7, suppose we apply $\widehat{\mathcal{T}}_{k,m}$ to $\widehat{Q}_{k,m}(s_{g},F_{z_{\Delta}},a_{g})$ and $\widehat{Q}_{k,m}^{\prime}(s_{g},F_{z_{\Delta}},a_{g})$. Then:
@@
$\|\widehat{\mathcal{T}}_{k,m}\widehat{Q}_{k}-\widehat{\mathcal{T}}_{k,m}\widehat{Q}^{\prime}_{k}\|_{\infty}$ $=\frac{\gamma}{m}\left\|\sum_{\ell\in[m]}\left(\max_{a^{\prime}\in\mathcal{A}}\widehat{Q}_{k}(s_{g}^{\ell},F_{z_{\Delta}^{\ell}},a_{g}^{\prime})-\max_{a^{\prime}\in\mathcal{A}}\widehat{Q}^{\prime}_{k}(s_{g}^{\ell},F_{z_{\Delta}^{\ell}},a_{g}^{\prime})\right)\right\|_{\infty}$
$\leq\gamma\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},s_{g}^{\prime}\in\mathcal{S}_{g},z_{\Delta}\in\mathcal{Z}_{l}^{k}\end{subarray}}|\widehat{Q}_{k}(s_{g}^{\prime},F_{z_{\Delta}^{\prime}},a_{g}^{\prime})-\widehat{Q}^{\prime}_{k}(s_{g}^{\prime},F_{z_{\Delta}^{\prime}},a_{g}^{\prime})|\leq\gamma\|\widehat{Q}_{k}-\widehat{Q}^{\prime}_{k}\|_{\infty}$
@@
The first inequality uses the triangle inequality and the general property $|\max_{a\in A}f(a)-\max_{b\in A}f(b)|\leq\max_{c\in A}|f(a)-f(b)|$. In the last line, we recover the definition of infinity norm.∎
Remark B.9.
@@
The $\gamma$-contractivity of $\widehat{\mathcal{T}}_{k}$ and $\widehat{\mathcal{T}}_{k,m}$ attracts the trajectory between two $\widehat{Q}_{k}$ and $\widehat{Q}_{k,m}$ functions on the same state-action tuple by $\gamma$ at each step. Repeatedly applying the Bellman operators produces a unique fixed-point from the Banach fixed-point theorem which we introduce in Definitions B.10 and B.11.
Definition B.10 ($\widehat{Q}_{k}^{*}$-function).
@@
Suppose $\widehat{Q}_{k}^{0}:=0$ and let $\widehat{Q}_{k}^{t+1}(s_{g},F_{z_{\Delta}},a_{g})=\widehat{\mathcal{T}}_{k}\widehat{Q}_{k}^{t}(s_{g},F_{z_{\Delta}},a_{g})$ for $t\in\mathbb{N}$. Denote the fixed-point of $\widehat{\mathcal{T}}_{k}$ by $\widehat{Q}^{*}_{k}$ such that $\widehat{\mathcal{T}}_{k}\widehat{Q}^{*}_{k}(s_{g},F_{z_{\Delta}},a_{g})=\widehat{Q}^{*}_{k}(s_{g},F_{z_{\Delta}},a_{g})$.
Definition B.11 ($\widehat{Q}_{k,m}^{\mathrm{est}}$-function).

### v3

@@ Acknowledgements.
We gratefully acknowledge insightful discussions with Siva Theja Maguluri, Yiheng Lin, Yi Wu, and Sarah Liaw.
@@ Appendix A Mathematical Background and Additional Remarks
  • Appendix B presents numerical simulations on the performance of SUBSAMPLE-MFQ on the Gaussian squeeze and constrained exploration tasks.
  • Appendix C presents notation and basic lemmas involving the learned $\hat{Q}_{k}$-function, as well as a stable and practical implementation of Algorithm 1.
  • Appendix D presents a proof sketch of our main result in Theorem 4.4.
  • Appendix E presents the proof of the Lipschitz continuity between $\hat{Q}_{k}$ and $Q^{*}$
  • Appendix F presents our bound on the TV-distance in Theorem D.2.
  • Appendix G proves the bound on the optimality gap between the learned policy $\tilde{\pi}_{k,m}^{\mathrm{est}}$ and the optimal policy $\pi^{*}$.
  • Appendix H presents an extension of the result to stochastic rewards.
  • Appendix I presents an extension of the result to off-policy learning that follows by using historical data.
  • Appendix J presents an extension of the result to continuous state/action spaces under a linear MDP assumption.
@@
$a_{1},\dots,a_{n}$ $a_{1},\dots,a_{n}\in\mathcal{A}_{l}^{n}$ are the actions of the local agents $1,\dots,n$;
$s_{1},\dots,s_{n}$ $s_{1},\dots,s_{n}\in\mathcal{S}_{l}^{n}$ are the states of the local agents $1,\dots,n$;
$a$ $a=(a_{g},a_{1},\dots,a_{n})\in\mathcal{A}_{g}\times\mathcal{A}_{l}^{n}$ is the tuple of actions of all agents;
$s$ $s=(s_{g},s_{1},\dots,s_{n})\in\mathcal{S}_{g}\times\mathcal{S}_{l}^{n}$ is the tuple of states of all agents;
$z_{i}$ $z_{i}=(s_{i},a_{i})\in\mathcal{Z}_{l}$, for $i\in[n]$;
$\mu_{k}(\mathcal{Z}_{l})$ $\mu_{k}(\mathcal{Z}_{l})=\{0,1/k,2/k,\dots,1\}^{|\mathcal{Z}_{l}|}$;
$\mu(\mathcal{Z}_{l})$ $\mu(\mathcal{Z}_{l})\coloneq\mu_{n}(\mathcal{Z}_{l})\{0,1/n,2/n,\dots,1\}^{|\mathcal{Z}_{l}|}$;
$s_{\Delta}$ For $\Delta\subseteq[n]$, and a collection of variables $\{s_{1},\dots,s_{n}\}$, $s_{\Delta}\coloneq\{s_{i}:i\in\Delta\}$;
@@
$\hat{\pi}_{k}^{*}$ $\hat{\pi}_{k}^{*}$ is the optimal deterministic policy function on a constrained system of
    $|\Delta|=k$ local agents;
$\tilde{\pi}^{\mathrm{est}}_{k}$ $\tilde{\pi}^{\mathrm{est}}_{k}$ is the stochastic policy map learned with parameter $k$ such that $a\sim\tilde{\pi}_{k}^{\mathrm{est}}(s)$;
$P_{g}(\cdot|s_{g},a_{g})$ $P_{g}(\cdot|s_{g},a_{g})$ is the stochastic transition kernel for the state of the global agent;
$P_{l}(\cdot|a_{i},s_{i},s_{g})$ $P_{l}(\cdot|a_{i},s_{i},s_{g})$ is the stochastic transition kernel for the state of local agent $i\in[n];$
$r_{g}(s_{g},a_{g})$ $r_{g}$ is the global agent’s component of the reward;
$r_{l}(s_{i},s_{g},a_{i})$ $r_{l}$ is the component of the reward for local agent $i\in[n]$;
$r(s,a)$ $r(s,a)=r_{g}(s_{g},a_{g})+\frac{1}{n}\sum_{i\in[n]}r_{l}(s_{i},s_{g},a_{i})$ is the reward of the system;
$r_{\Delta}(s,a)$ $r_{\Delta}(s,a)=r_{g}(s_{g},a_{g})+\frac{1}{|\Delta|}\sum_{i\in\Delta}r_{l}(s_{i},s_{g},a_{i})$ is the constrained system’s reward
    with $|\Delta|=k$ local agents;
$\mathcal{T}$ $\mathcal{T}$ is the centralized Bellman operator;
$\hat{\mathcal{T}}_{k}$ $\hat{\mathcal{T}}_{k}$ is the Bellman operator on a constrained system of $|\Delta|=k$ local agents;
@@
Given metric spaces $(\mathcal{X},d_{\mathcal{X}})$ and $(\mathcal{Y},d_{\mathcal{Y}})$ and a constant $L>0$, a map $f:\mathcal{X}\to\mathcal{Y}$ is $L$-Lipschitz continuous if for all $x,y\in\mathcal{X}$, $d_{\mathcal{Y}}(f(x),f(y))\leq L\cdot d_{\mathcal{X}}(x,y)$.
Theorem A.2 (Banach-Caccioppoli fixed point theorem [Banach, 1922]).
@@
Appendix B Numerical Experiments
This section provides numerical simulations for the examples outlined in Section 2. All experiments were run on a 2-core CPU server with 12GB RAM. We chose a parameter complexity for each simulation that was sufficient to emphasize characteristics of the theory, such as the complexity improvement and the decaying optimality gap.
Figure 2: a) Reward optimality gap (log scale) with ${\pi}_{k,m}^{\mathrm{est}}$ running $300$ iterations. b) Computation time (in minutes) against sampling parameter $k$, for $k\leq n=8$, to learn policy $\hat{\pi}_{k,m}^{\mathrm{est}}$. c) Discounted cumulative rewards for $k\leq n=50$.
B.1 Constrained Exploration
[…]
Intuitively, the reward function is designed to force the global agent to oscillate vertically in the grid, while forcing the local agents to oscillate horizontally around the global agent, thereby simulating a constrained exploration. This model has been studied previously in Lin et al. [2023b].
For this task, we ran a simulation with $n=8$ agents, with $m=20$ samples in the empirically adapted Bellman operator. We provide simulation results in Figure 2a. We observe monotonic improvements in the cumulative discounted rewards as $k\to n$. Since $k=n$ recovers value-iteration and mean-field MARL algorithms, the reward at $k=n$ is the baseline we compare our algorithm to. When $k<n$, we observe that the reward accrued by SUBSAMPLE-MFQ is only marginally less than the reward gained by value-iteration.
B.2 Gaussian Squeeze (GS)
In this task, $n$ homogeneous agents determine their individual action $a_{i}$ to jointly maximize the objective $r(x)=xe^{-(x-\mu)^{2}/\sigma^{2}}$, where $x=\sum_{i=1}^{n}a_{i}$, $a_{i}\in\{0,\dots,9\}$, and $\mu$ and $\sigma$ are the pre-defined mean and variance of the system. We consider a homogeneous system devoid of a global agent to match the mean-field setting in Yang et al. [2018], where the richness of our model setting can still express GS. We use the global agent to model the state of the system, given by the number of vehicles in each controller’s lane.
We set $\mathcal{S}_{l}=[4]$, $\mathcal{A}_{l}=\{0,\dots,4\}$, and $\mathcal{S}_{g}=\mathcal{S}_{l}$ such that $s_{g}=\left\lceil\frac{1}{n}\sum_{i=1}^{n}s_{i}\right\rceil$, and $\mathcal{A}_{g}=\{0\}$. The transition functions are given by $s_{i}(t+1)=s_{i}(t)-1\cdot\mathbbm{1}\{s_{i}(t)>s_{g}(t)\}+\mathrm{Ber}(p)$ and $s_{g}(t+1)=\left\lceil\frac{1}{n}\sum_{i=1}^{n}s_{i}(t+1)\right\rceil$, where $\mathrm{Ber}(p)$ is a Bernoulli random variable with parameter $p>0$. Finally, the global agent’s reward function is given by $r_{g}(s_{g},a_{g})=-s_{g}$ and the local agent’s reward function is given by $r_{l}(s_{i},s_{g},a_{i})=s_{i}\cdot e^{-(s_{i}-s_{g})^{2}/4}-\max\{s_{i},a_{i}\}$.
For this task, we ran a small-scale simulation with $n=8$ agents, and a large-scale simulation with $n=50$ agents, and used $m=20$ samples in the empirical Bellman operator. We provide simulation results in Figure 2b and Figure 2c, where Figure 2b demonstrates the exponential improvement in computational complexity of SUBSAMPLE-MFQ, and Figure 2c demonstrates a monotonic improvement in the cumulative rewards, consistent with Theorem 4.4. Here, both metrics outperform the mean-field value iteration benchmark.
Appendix C Notation and Basic Lemmas
@@
Definition C.1 (Bellman Operator $\mathcal{T}$).
@@
Definition C.2 (Adapted Bellman Operator $\hat{\mathcal{T}}_{k}$).
The adapted Bellman operator updates a smaller $Q$ function (which we denote by $\hat{Q}_{k}$), for a surrogate system with the global agent and $k\in[n]$ local agents denoted by $\Delta$, using mean-field value iteration and $j\in\Delta$ such that:
$\hat{\mathcal{T}}_{k}\hat{Q}_{k}^{t}(s_{g},s_{j},F_{z_{\Delta\setminus j}},a_{j},a_{g}):=r_{\Delta}(s,a)+\gamma\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim P_{g}(\cdot|s_{g},a_{g}),\\ s_{i}^{\prime}\sim P_{l}(\cdot|s_{i},s_{g},a_{i}),\forall i\in\Delta\end{subarray}}\max_{a^{\prime}\in\mathcal{A}}\hat{Q}_{k}^{t}(s_{g}^{\prime},s_{j}^{\prime},F_{z^{\prime}_{\Delta\setminus j}},a^{\prime}_{j},a^{\prime}_{g})$ (14)
Definition C.3 (Empirical Adapted Bellman Operator $\hat{\mathcal{T}}_{k,m}$).
The empirical adapted Bellman operator $\hat{\mathcal{T}}_{k,m}$ empirically estimates the adapted Bellman operator update using mean-field value iteration by drawing $m$ random samples of $s_{g}\sim P_{g}(\cdot|s_{g},a_{g})$ and $s_{i}\sim P_{l}(\cdot|s_{i},s_{g},a_{i})$ for $i\in\Delta$, where for $\ell\in[m]$, the $\ell$’th random sample is given by $s_{g}^{\ell}$ and $s_{\Delta}^{\ell}$, and $j\in\Delta$:
$\hat{\mathcal{T}}_{k,m}\hat{Q}_{k,m}^{t}(s_{g},s_{j},F_{z_{\Delta\setminus j}},a_{j},a_{g}):=r_{\Delta}(s,a)+\frac{\gamma}{m}\sum_{\ell\in[m]}\max_{a^{\prime}\in\mathcal{A}}\hat{Q}_{k,m}^{t}(s_{g}^{\ell},s_{j}^{\ell},F_{z_{\Delta\setminus j}^{\ell}},a_{j}^{\ell},a_{g}^{\ell})$ (15)
Lemma C.4.
For any $\Delta\subseteq[n]$ such that $|\Delta|=k$, suppose $0\leq r_{\Delta}(s,a)\leq\tilde{r}$. Then, for all $t\in\mathbb{N}$, $\hat{Q}_{k}^{t}\leq\frac{\tilde{r}}{1-\gamma}$.
@@
The proof follows by induction on $t$. The base case follows from $\hat{Q}_{k}^{0}:=0$. For the induction, note that by the triangle inequality $\|\hat{Q}_{k}^{t+1}\|_{\infty}\leq\|r_{\Delta}\|_{\infty}+\gamma\|\hat{Q}_{k}^{t}\|_{\infty}\leq\tilde{r}+\gamma\frac{\tilde{r}}{1-\gamma}=\frac{\tilde{r}}{1-\gamma}$. ∎
Remark C.5.
By the law of large numbers, $\lim_{m\to\infty}\hat{\mathcal{T}}_{k,m}=\hat{\mathcal{T}}_{k}$, where the error decays in $O(1/\sqrt{m})$ by the Chernoff bound. Also, $\hat{\mathcal{T}}_{n}\coloneq\mathcal{T}$. Further, Lemma C.4 is independent of the choice of $k$. Therefore, for $k=n$, this implies an identical bound on $Q^{t}$. An identical argument implies the same bound on $\hat{Q}_{k,m}^{t}$.
$\mathcal{T}$ satisfies a $\gamma$-contractive property under the infinity norm [Watkins and Dayan, 1992]. We similarly show that $\hat{\mathcal{T}}_{k}$ and $\hat{\mathcal{T}}_{k,m}$ satisfy a $\gamma$-contractive property under infinity norm in Lemmas C.6 and C.7.
Lemma C.6.
$\hat{\mathcal{T}}_{k}$ satisfies the $\gamma$-contractive property under infinity norm:
$\|\hat{\mathcal{T}}_{k}\hat{Q}_{k}^{\prime}-\hat{\mathcal{T}}_{k}\hat{Q}_{k}\|_{\infty}\leq\gamma\|\hat{Q}_{k}^{\prime}-\hat{Q}_{k}\|_{\infty}$ (16)
@@
Suppose we apply $\hat{\mathcal{T}}_{k}$ to $\hat{Q}_{k}(s_{g},F_{z_{\Delta}},a_{g})$ and $\hat{Q}^{\prime}_{k}(s_{g},F_{z_{\Delta}},a_{g})$ for $|\Delta|=k$. Then:
$\|\hat{\mathcal{T}}_{k}\hat{Q}_{k}^{\prime}-\hat{\mathcal{T}}_{k}\hat{Q}_{k}\|_{\infty}$
$=\gamma\max_{\begin{subarray}{c}s_{g}\in\mathcal{S}_{g},\\ a_{g}\in\mathcal{A}_{g},\\ F_{z_{\Delta}}\in\mu_{k}{(\mathcal{Z}_{l})}\end{subarray}}\!\left|\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim P_{g}(\cdot|s_{g},a_{g}),\\ s_{i}^{\prime}\sim P_{l}(\cdot|s_{i},s_{g},a_{i}),\\ \forall i\in\Delta,\\ \end{subarray}}\max_{a^{\prime}\in\mathcal{A}}\hat{Q}_{k}^{\prime}(s_{g}^{\prime},F_{z_{\Delta}^{\prime}},a_{g}^{\prime})-\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim P_{g}(\cdot|s_{g},a_{g}),\\ s_{i}^{\prime}\sim P_{l}(\cdot|s_{i},s_{g},a_{i}),\\ \forall i\in\Delta^{\prime}\end{subarray}}\max_{a^{\prime}\in\mathcal{A}}\hat{Q}_{k}(s_{g}^{\prime},F_{z_{\Delta}^{\prime}},a_{g}^{\prime})\right|$
$\leq\gamma\max_{\begin{subarray}{c}s_{g}^{\prime}\in\mathcal{S}_{g},F_{z_{\Delta}}\in\mu_{k}{(\mathcal{Z}_{l})},a^{\prime}\in\mathcal{A}\end{subarray}}\left|\hat{Q}_{k}^{\prime}(s_{g}^{\prime},F_{z_{\Delta}^{\prime}},a_{g}^{\prime})-\hat{Q}_{k}(s_{g}^{\prime},F_{z_{\Delta}^{\prime}},a_{g}^{\prime})\right|=\gamma\|\hat{Q}_{k}^{\prime}-\hat{Q}_{k}\|_{\infty}$
@@
Lemma C.7.
$\hat{\mathcal{T}}_{k,m}$ satisfies the $\gamma$-contractive property under infinity norm.
@@
Similarly to Lemma C.6, suppose we apply $\hat{\mathcal{T}}_{k,m}$ to $\hat{Q}_{k,m}(s_{g},F_{z_{\Delta}},a_{g})$ and $\hat{Q}_{k,m}^{\prime}(s_{g},F_{z_{\Delta}},a_{g})$. Then:
@@
$\|\hat{\mathcal{T}}_{k,m}\hat{Q}_{k}-\hat{\mathcal{T}}_{k,m}\hat{Q}^{\prime}_{k}\|_{\infty}$ $=\frac{\gamma}{m}\left\|\sum_{\ell\in[m]}\left(\max_{a^{\prime}\in\mathcal{A}}\hat{Q}_{k}(s_{g}^{\ell},F_{z_{\Delta}^{\ell}},a_{g}^{\prime})-\max_{a^{\prime}\in\mathcal{A}}\hat{Q}^{\prime}_{k}(s_{g}^{\ell},F_{z_{\Delta}^{\ell}},a_{g}^{\prime})\right)\right\|_{\infty}$
$\leq\gamma\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},s_{g}^{\prime}\in\mathcal{S}_{g},z_{\Delta}\in\mathcal{Z}_{l}^{k}\end{subarray}}|\hat{Q}_{k}(s_{g}^{\prime},F_{z_{\Delta}^{\prime}},a_{g}^{\prime})-\hat{Q}^{\prime}_{k}(s_{g}^{\prime},F_{z_{\Delta}^{\prime}},a_{g}^{\prime})|$
$=\gamma\|\hat{Q}_{k}-\hat{Q}^{\prime}_{k}\|_{\infty}$
@@
The first inequality uses the triangle inequality and the general property $|\max_{a\in A}f(a)-\max_{b\in A}f(b)|\leq\max_{c\in A}|f(a)-f(b)|$. The last line recovers the definition of infinity norm.∎
Remark C.8.
@@
The $\gamma$-contractivity of $\hat{\mathcal{T}}_{k}$ and $\hat{\mathcal{T}}_{k,m}$ attracts the trajectory between two $\hat{Q}_{k}$ and $\hat{Q}_{k,m}$ functions on the same state-action tuple by $\gamma$ at each step. Repeatedly applying the Bellman operators produces a unique fixed-point from the Banach fixed-point theorem which we introduce in Definitions C.9 and C.10.
Definition C.9 ($\hat{Q}_{k}^{*}$-function).
@@
Suppose $\hat{Q}_{k}^{0}:=0$ and let $\hat{Q}_{k}^{t+1}(s_{g},F_{z_{\Delta}},a_{g})=\hat{\mathcal{T}}_{k}\hat{Q}_{k}^{t}(s_{g},F_{z_{\Delta}},a_{g})$ for $t\in\mathbb{N}$. Denote the fixed-point of $\hat{\mathcal{T}}_{k}$ by $\hat{Q}^{*}_{k}$ such that $\hat{\mathcal{T}}_{k}\hat{Q}^{*}_{k}(s_{g},F_{z_{\Delta}},a_{g})=\hat{Q}^{*}_{k}(s_{g},F_{z_{\Delta}},a_{g})$.
Definition C.10 ($\hat{Q}_{k,m}^{\mathrm{est}}$-function).